# Distributed Stochastic Gradient Descent with Staleness: A Stochastic Delay Differential Equation Based Framework

### v1

@@ II-A Asynchronous Distributed Optimization
in which $\nabla f(\boldsymbol{x})$ is the gradient of the objective function $f(\cdot)$ at $\boldsymbol{x}$, $\tau(t)$ is the stochastic process characterizing the gradient staleness in the SDDE, $\boldsymbol{B}(t)$ or $B(t)$ is an $N$-dimensional or one-dimensional standard Brownian motion or Wiener process, and $\sigma>0$ is the standard variance of the Brownian motion. [^2] To be rigorous, given the learning rate $\eta$ and the standard variance of $\boldsymbol{w}[\cdot]$, denoted by $\breve{\sigma}$, the relationship between the standard variance of Brownian motion in (4) and the Gaussian noise in (2) is $\sigma=\frac{1}{\sqrt{\eta}}\breve{\sigma}$.
@@ II-B Event-Triggered Distributed Optimization
In this subsection, we present the implementation of ET-SDG. In distributed SGD with limited bandwidth, event-triggered SGD can be adopted to effectively reduce bandwidth usage. In particular, different from asynchronous SGD, time is partitioned into fixed-length time slots in the ET-SDG setup. Let $T_{s}$ denote the time slot length. The central server performs updates of the model at wall-clock time $jT_{s}$ for $j\in\mathbb{N}^{+}$. Further, the communication between the workers and the central server follows an event-triggered manner. In this scheme, each worker communicates its gradient to the server only upon a certain event is triggered. More specifically, the triggering condition is that the difference between the current gradient and the last updated gradient is greater than a pre-given threshold.
@@
In particular, the central server broadcasts the current model to all the workers at each iteration. Based on the received model and the local dataset, each worker computes a gradient. Further, it is assumed that each worker could finish the computation of the gradient within a timeslot. More specifically, the gradient computed by Worker $\kappa$ at iteration $j$ is denoted by ${\boldsymbol{g}}^{\kappa}[j]=\nabla f(\breve{\boldsymbol{x}}[j])+\boldsymbol{w}^{\kappa}[j]$. Note that the workers skip transmission of the current gradient at the epoch when the current gradient and the last updated gradient do not differ too much [24]. More specifically, the triggering criterion is given by $\left\|{\boldsymbol{g}}^{\kappa}[j]-\hat{\boldsymbol{g}}^{\kappa}[j^{-}]\right\|_{2}\geq\xi$ in which ${\boldsymbol{g}}^{\kappa}[j]$ is the newly computed gradient, $\hat{\boldsymbol{g}}^{\kappa}[j^{-}]$ is Worker $\kappa$’s last updated gradient before time $j^{-}$, $\left\|\cdot\right\|_{2}$ denotes the 2-norm, and $\xi$ is the threshold in event-triggered SGD. Further, let $\hat{\boldsymbol{g}}^{\kappa}[j]={\boldsymbol{g}}^{\kappa}[j-\breve{\tau}^{\kappa}[j]],$ in which $\breve{\tau}^{\kappa}[j]$ is the gradient staleness arises in event-triggered SGD and $\breve{\tau}^{\kappa}[j]=0$ if Worker $\kappa$ update its gradient at time $j$. In this way, the update rule of the global model in event-triggered SGD in discrete-time is $\breve{\boldsymbol{x}}[j+1]=\breve{\boldsymbol{x}}[j]-\frac{\eta}{K}\sum_{\kappa=1}^{K}\hat{\boldsymbol{g}}^{\kappa}[j-\breve{\tau}^{\kappa}[j]].$ Slightly different from the asynchronous SGD setting, the learning rate is $\eta/K$ with $K$ workers in event-triggered SGD. Correspondingly, the update rule of event-triggered SGD can be also be viewed as the discrete-time version of the continuous-time process (4), in which $\tau(t)=\eta\breve{\tau}^{\kappa}[j]$ for $t\in((j-1)\eta+(\kappa-1)\eta/K,(j-1)\eta+\kappa\eta/K)$ and $\breve{\boldsymbol{x}}[j]\approx\boldsymbol{x}(j\eta)$ in event-triggered SGD.
@@
In Section III, we aim to reveal the relationship between gradient staleness and noise and the convergence rate of the SDDE. In Section IV, we shall reveal how the protocol parameters $K$, $B$, and the computational/communication delays determine $N_{a}(t)$ and the step staleness $\breve{\tau}$, as well as the gradient dropout schemes that can be adopted to trade update frequency for a smaller step staleness. According to the SDDE approximation, we bridge the convergence rate of asynchronous SGD and the learning rate, the objective function, and the protocol parameters, based on which the protocol design criteria are further revealed.
@@ III A Unified Framework for Convergence Analysis of Distributed SGD
In this section, we present theoretical performance analyses for the relationship between the gradient staleness and noise and the convergence of distributed SGD through an SDDE-based approach. To obtain analytical results and find more insights, we first focus on the solvable cases, as was done in, e.g., [38, 37]. Here, we assume that the objective function can be locally approximated by a quadratic function [46, 47, 48]. Specifically, the objective function takes the form $f(\boldsymbol{x})=\frac{1}{2}(\boldsymbol{x}-\boldsymbol{x}^{*})^{T}\boldsymbol{V}(\boldsymbol{x}-\boldsymbol{x}^{*}),$ in which $\boldsymbol{V}\in\mathbb{R}^{N\times N}$ is a positive-define matrix, $\boldsymbol{x}\in\mathbb{R}^{N}$ is the objective variable, $\boldsymbol{x}^{*}\in\mathbb{R}^{N}$ is a constants, and $[\cdot]^{T}$ denotes transpose.
@@ III-A A Unified Framework Based on SDDE
For the sake of discussion, ${\boldsymbol{V}}$ is assumed to be a diagonal matrix, i.e., $\boldsymbol{V}=\text{diag}(v_{1},v_{2},...,v_{N})$ and it is assumed that $\boldsymbol{x}^{*}=\boldsymbol{0}$ in which $\boldsymbol{0}$ is the vector of zeros. [^4] The assumption that ${\boldsymbol{V}}$ is a diagonal matrix is with loss of generality. In case $\boldsymbol{V}$ is not a diagonal matrix, it holds that $\boldsymbol{V}=\boldsymbol{U}\tilde{\boldsymbol{V}}\boldsymbol{U}^{T}$ in which $\tilde{\boldsymbol{V}}$ is a diagonal matrix by eigenvalue decomposition. By defining $\tilde{\boldsymbol{x}}=\boldsymbol{U}^{T}(\boldsymbol{x}-\boldsymbol{x}^{*})$, we have $f\left(\tilde{\boldsymbol{x}}\right)=\tilde{\boldsymbol{x}}^{T}\tilde{\boldsymbol{V}}\tilde{\boldsymbol{x}}$. By this means, we focus on the diagonal $\boldsymbol{V}$ in the subsequent discussion. With a quadratic objective function, the continuous-time approximation of distributed SGD with gradient staleness (4) can be represented as:
$\mathrm{d}{\boldsymbol{x}}(t)=-{\boldsymbol{V}}{\boldsymbol{x}}(t-\tau(t))\mathrm{d}t+\sigma\mathrm{d}\boldsymbol{B}(t),$ (5)
@@
$\mathrm{d}x_{i}(t)=-v_{i}x_{i}(t-\tau(t))\mathrm{d}t+\sigma\mathrm{d}B(t),$ (6)
for $i=1,2,...,N.$ When $\sigma=0$, implying that each gradient is computed based on the entire training dataset where the gradient noise is negligible, Eq. (6) reduces to the deterministic differential equation given by $\mathrm{d}x_{i}(t)=-v_{i}x_{i}(t-\tau(t))\mathrm{d}t$. In addition, when $\tau(t)=0$, which corresponds to synchronous SGD, we obtain $\mathrm{d}x_{i}(t)=-v_{i}x_{i}(t)\mathrm{d}t+\sigma\mathrm{d}B(t)$ which is an Ornstein-Uhlenbeck process, abbreviated as OU-process. Also, it is observed that the OU process with delay (5) along each eigenvector direction is independent, allowing us to first focus on the one-dimensional case (6) then extend our discussion to the multi-dimensional case.
Next, we demonstrate that the impact of the gradient noise and the gradient staleness can be separately considered taking the case that $\tau$ is a constant as an example. [^5] Note that the gradient staleness $\tau$ in SDDE is approximately a constant the variance of computation delay of each worker is small with a large worker number in asynchronous SGD, as will be demonstrated in Section IV. In this case, the analysis of the behavior of SGD can be based on viewing SGD as a discretization of the following associated stochastic delay differential equation, i.e.,
@@
Let $\tilde{x}_{d}(t)$ denote a solution to Eq. (8). Then, the solution to Eq. (7) can be given by $x_{s}(t)=\tilde{x}_{d}(t)+\sigma\int_{0}^{t}x_{0}(t-s)\mathrm{d}B(t),$ in which $x_{0}(t)=\sum_{k=0}^{\lfloor\frac{t}{\tau}\rfloor}\left[\frac{{v}^{k}}{k!}(t-k\tau)\right]$ in which $t\in(0,\infty)$. In addition, we have $\tilde{x}_{d}(t)=x_{0}(t)\varphi(0)+{v}\int_{-\tau}^{0}x_{0}(t-s-\tau)\varphi(s)\mathrm{d}s$. It can be noticed that the difference between $x_{s}(t)$ and $\tilde{x}_{d}(t)$ is only given by a Gaussian distributed random variable, the variance of which is joint determined by the gradient staleness $\tau$ and the standard variance of the Brownian motion $\sigma$.
In the following discussions, we shall present performance analyses for the convergence rate of distributed SGD given gradient noise or gradient staleness through the SDDE-based approach.

### v2

@@ II-A Asynchronous Distributed Optimization
in which $\nabla f(\boldsymbol{x})$ is the gradient of the objective function $f(\cdot)$ at $\boldsymbol{x}$, $\tau(t)$ is the stochastic process characterizing the gradient staleness in the SDDE, $\boldsymbol{B}(t)$ or $B(t)$ is an $N$-dimensional or one-dimensional standard Brownian motion or Wiener process, and $\sigma>0$ is the standard variance of the Wiener process, in which a standard $N$-dimensional Wiener process is a vector-valued stochastic process $\boldsymbol{B}(t)=\left(B_{1}(t),B_{2}(t),\ldots,B_{N}(t)\right)$, where each $B_{i}(t)$ is an independent standard one-dimensional Wiener process for $i=1,2,\ldots,N$ [^3] To be rigorous, given the learning rate $\eta$ and the standard variance of $\boldsymbol{w}[\cdot]$, denoted by $\breve{\sigma}$, the relationship between the standard variance of Wiener process in (5) and the Gaussian noise in (3) is $\sigma=\frac{1}{\sqrt{\eta}}\breve{\sigma}$.
@@
In Section III, we aim to reveal the relationship between gradient staleness and noise and the convergence rate of the SDDE. In Section IV, we shall reveal how the protocol parameters $K$, $B$, and the computational/communication delays determine $N_{a}(t)$ and the step staleness $\breve{\tau}$, as well as the gradient dropout schemes that can be adopted to trade update frequency for a smaller step staleness. According to the SDDE approximation, we bridge the convergence rate of ASGD and the learning rate, the objective function, and the protocol parameters, based on which the protocol design criteria are further revealed.
@@ II-B Event-Triggered Distributed Optimization
In this subsection, we present the implementation of ET-SDG. In distributed SGD with limited bandwidth, event-triggered SGD can be adopted to effectively reduce bandwidth usage. In particular, different from ASGD, time is partitioned into fixed-length time slots in the ET-SDG setup. Let $T_{s}$ denote the time slot length. The central server performs updates of the model at wall-clock time $jT_{s}$ for $j\in\mathbb{N}^{+}$. Further, the communication between the workers and the central server follows an event-triggered manner. In this scheme, each worker communicates its gradient to the server only upon a certain event is triggered. More specifically, the triggering condition is that the difference between the current gradient and the last updated gradient is greater than a pre-given threshold.
@@
In particular, the central server broadcasts the current model to all the workers at each iteration. Based on the received model and the local dataset, each worker computes a gradient. Further, it is assumed that each worker could finish the computation of the gradient within a timeslot. More specifically, the gradient computed by Worker $\kappa$ at iteration $j$ is denoted by ${\boldsymbol{g}}^{\kappa}[j]=\nabla f(\breve{\boldsymbol{x}}[j])+\boldsymbol{w}^{\kappa}[j]$. Note that the workers skip transmission of the current gradient at the epoch when the current gradient and the last updated gradient do not differ too much [25]. More specifically, the triggering criterion is given by $\left\|{\boldsymbol{g}}^{\kappa}[j]-\hat{\boldsymbol{g}}^{\kappa}[j^{-}]\right\|_{2}\geq\xi$ in which ${\boldsymbol{g}}^{\kappa}[j]$ is the newly computed gradient, $\hat{\boldsymbol{g}}^{\kappa}[j^{-}]$ is Worker $\kappa$’s last updated gradient before time $j^{-}$, $\left\|\cdot\right\|_{2}$ denotes the 2-norm, and $\xi$ is the threshold in event-triggered SGD. Further, let $\hat{\boldsymbol{g}}^{\kappa}[j]={\boldsymbol{g}}^{\kappa}[j-\breve{\tau}^{\kappa}[j]],$ in which $\breve{\tau}^{\kappa}[j]$ is the gradient staleness that arises in event-triggered SGD and $\breve{\tau}^{\kappa}[j]=0$ if Worker $\kappa$ pushes its gradient at time $j$. In this way, the update rule of the global model in event-triggered SGD in discrete-time is $\breve{\boldsymbol{x}}[j+1]=\breve{\boldsymbol{x}}[j]-\frac{\eta}{K}\sum_{\kappa=1}^{K}\hat{\boldsymbol{g}}^{\kappa}[j-\breve{\tau}^{\kappa}[j]].$ Slightly different from the ASGD setting, the learning rate is $\eta/K$ with $K$ workers in event-triggered SGD. Correspondingly, the update rule of event-triggered SGD can also be viewed as the discrete-time version of the continuous-time process (5), in which $\tau(t)=\eta\breve{\tau}^{\kappa}[j]$ for $t\in((j-1)\eta+(\kappa-1)\eta/K,(j-1)\eta+\kappa\eta/K)$ and $\breve{\boldsymbol{x}}[j]\approx\boldsymbol{x}(j\eta)$ in event-triggered SGD.
@@
In Section III, we aim to reveal the relationship between gradient staleness and noise and the convergence rate of the SDDE. In Section IV, we shall reveal how the protocol parameters $K$, $B$, and the computational/communication delays determine $N_{a}(t)$ and the step staleness $\breve{\tau}$, as well as the gradient dropout schemes that can be adopted to trade update frequency for a smaller step staleness. According to the SDDE approximation, we bridge the convergence rate of ASGD and the learning rate, the objective function, and the protocol parameters, based on which the protocol design criteria are further revealed.
II-C Definitions of the Key Mathematical Concepts
In this subsection, we present definitions of the key mathematical concepts of this paper. Let us first introduce the concept of the OU process.
Definition 1
A one-dimensional OU process $O(t)$ is the solution to the stochastic differential equation
${\rm d}O(t)=\alpha(\beta-O(t)){\rm d}t+\sigma{\rm d}B(t),\text{ for }t>0,$ (6)
where $\alpha$, $\beta$, and $\sigma\geq 0$ are real constants, and $B(t)$ is a standard Wiener process.
A one-dimensional OU process $O(t)$ can also be defined in terms of a stochastic integral as
$O(t)=\beta(1-\exp(-\alpha t))+$ (7)
$\sigma\exp(-\alpha t)\int_{0}^{t}\exp(\alpha s){\rm d}B(s)+O(0)\exp(-\alpha t).$ (8)
where $O(0)$ is a random variable denoting the initial state. Further, conditional on $O(0)$ where $\mathbb{E}\{O(0)\}<\infty$, $O(t)$ is a Gaussian random variable with mean $\mathbb{E}\{O(t)\}=\beta(1-\exp(-\alpha t))+\exp(-\alpha t)\mathbb{E}\{O(0)\}$. In this paper, our main focus is on the case that $\beta=0$. In this case, it holds that
$O(t)=\sigma\exp(-\alpha t)\int_{0}^{t}\exp(\alpha s){\rm d}B(s)+O(0)\exp(-\alpha t).$ (9)
which is a Gaussian random variable with mean $\mathbb{E}\{O(t)\}=\exp(-\alpha t)\mathbb{E}\{O(0)\}$ and variance $\text{Var}\{O(t)\}=\frac{\sigma^{2}}{2\alpha}(1-\exp(-2\alpha t))$.
The stopping criterion of SGD is usually of the form $\|\nabla f(x)\|_{2}\leq\delta$, where $\delta$ is small and positive [48]. In this case, the training time of SGD can be characterized by the first hitting time of the objective variable to the set $S=\{x:\|\nabla f(x)\|_{2}\leq\delta\}$. In the following, let us present the definition of the first hitting time.
Definition 2
Let $\{X(t),{t\geq 0}\}$ denote a stochastic process in an $N$-dimensional spatial domain $\Omega\subseteq\mathbb{R}^{N}$, the first hitting time to a target set $\Omega_{t}\subset\Omega$ is defined as
$H:=\inf\{t\geq 0:X(t)\in\Omega_{t}\}.$ (10)
For Brownian motion with drift, we have the following lemma according to [49].
Lemma 1
The expectation of the first hitting time $H=\inf\{t>0:x_{1}\leq x(t)\leq x_{1}+\mathrm{d}x,x(0)=x_{0}\}$ of the one-dimensional stochastic process $\mathrm{d}x(t)=s(x)\mathrm{d}t+\sigma(x)\mathrm{d}B(t)$ is $\mathbb{E}\{H\}=\int_{x_{0}}^{x_{1}}\frac{1}{s(x)}\mathrm{d}x,$ in which $s(x)$ is the drift velocity.
For example, the fact that the expected hitting time of a given level by a standard Wiener process is infinite is a direct corollary of Lemma 1. For convenience, we summarize the definitions of important variables in Table I.
@@ III A Unified Framework for Convergence Analysis of Distributed SGD
In this section, we present theoretical performance analyses for the relationship between gradient staleness and noise and the convergence of distributed SGD through an SDDE-based approach. To obtain analytical results and find more insights, we focus on the solvable cases in this section, as was done in, e.g., [43, 44]. More specifically, in this section, our analysis focuses on the quadratic model [50, 51, 52]. Specifically, the objective function takes the form $f(\boldsymbol{x})=\frac{1}{2}(\boldsymbol{x}-\boldsymbol{x}^{*})^{T}\boldsymbol{V}(\boldsymbol{x}-\boldsymbol{x}^{*})+\boldsymbol{q}^{T}(\boldsymbol{x}-\boldsymbol{x}^{*})+\boldsymbol{c},$ in which $\boldsymbol{V}\in\mathbb{R}^{N\times N}$ is a symmetric matrix, $\boldsymbol{x}\in\mathbb{R}^{N}$ is the objective variable, $\boldsymbol{q},\boldsymbol{c}$, and $\boldsymbol{x}^{*}\in\mathbb{R}^{N}$ are constants, and $[\cdot]^{T}$ denotes transpose. In the literature, including [43, 44, 53], which uses differential equations to analyze optimization methods for deeper understanding, quadratic models are frequently employed to derive closed-form solutions.
@@ III-A A Unified Framework Based on SDDE
In this subsection, we first present the following discussion on the objective function. Note that $\boldsymbol{V}$ is a symmetric matrix. Therefore, it holds that $\boldsymbol{V}=\boldsymbol{U}\tilde{\boldsymbol{V}}\boldsymbol{U}^{T}$ in which $\tilde{\boldsymbol{V}}$ is a diagonal matrix and $\boldsymbol{U}$ is a unitary matrix by eigendecomposition. Hence, by defining $\tilde{\boldsymbol{x}}=\boldsymbol{U}^{T}(\boldsymbol{x}-\boldsymbol{x}^{*})$ and $\tilde{q}=\boldsymbol{q}^{T}\boldsymbol{U}$, the objective function can be rewritten as $f\left(\tilde{\boldsymbol{x}}\right)=\frac{1}{2}\tilde{\boldsymbol{x}}^{T}\tilde{\boldsymbol{V}}\tilde{\boldsymbol{x}}+\tilde{\boldsymbol{q}}^{T}\tilde{\boldsymbol{x}}+\boldsymbol{c}$. where $\tilde{\boldsymbol{V}}$ is a diagonal matrix. By omitting the tildes for brevity, the objective function can be formulated as $f(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{V}\boldsymbol{x}+\boldsymbol{q}^{T}\boldsymbol{x}+\boldsymbol{c}$. By this means, we assume the symmetric matrix $\boldsymbol{V}$ is a diagonal matrix given by $\boldsymbol{V}=\text{diag}(v_{1},v_{2},...,v_{N})$. without loss of generality. Further, the continuous-time approximation of distributed SGD with gradient staleness (5) can be represented as:
$\mathrm{d}{\boldsymbol{x}}(t)=-({{\boldsymbol{V}}{\boldsymbol{x}}(t-\tau(t))+\boldsymbol{q}})\mathrm{d}t+\sigma\mathrm{d}\boldsymbol{B}(t),$ (11)
@@
$\mathrm{d}x_{i}(t)=-(v_{i}x_{i}(t-\tau(t))+q_{i})\mathrm{d}t+\sigma\mathrm{d}B(t),$ (12)
for $i=1,2,...,N$, where $q_{i}$ is the $i$th element of $\boldsymbol{q}$. Further, by introducing $\bar{x}_{i}(t)=x_{i}(t)+\frac{q_{i}}{v_{i}}$ when $v_{i}\neq 0$, Eq. (12) can be rewritten as
$\mathrm{d}\bar{x}_{i}(t)=-v_{i}\bar{x}_{i}(t-\tau(t))\mathrm{d}t+\sigma\mathrm{d}B(t),\text{ for }v_{i}\neq 0.$ (13)
Therefore, for the sake of discussion, we make the assumption that $q_{i}=0$ for any $i$ when $v_{i}\neq 0$ without any loss of generality.
Remark:  Similar to most existing work in the literature on the continuous approximation analysis of SGD [9, 54, 44, 43], we focus on quadratic objective functions, as they allow for closed-form analysis of training time, convergence rate, and the stationary distribution. For example, in [55, Section 7.1.1], it is noted that insights into the optimization problem and the various techniques for solving it can be gained by considering a local quadratic approximation to the error function.
While this method can be extended to twice-differentiable functions $f(\boldsymbol{x})$ to analyze the local behavior of SGD, we focus on approximating $f(\boldsymbol{x})$ around a specific point $\boldsymbol{x}_{0}$ within its domain. For small neighborhoods around $\boldsymbol{x}_{0}$, we approximate $f(\boldsymbol{x})$ using a quadratic model function derived from Taylor expansion, as outlined by Assumption 1. For example, the discussion on the quadratic model is particularly relevant when SGD approaches a local or global minimum. In this case, $\boldsymbol{V}$ is a positive-defined matrix, meaning that all of its eigenvalues are greater than zero, and it holds that $\boldsymbol{q}=0$. On the other hand, when some eigenvalues of $\boldsymbol{V}$ are negative, this corresponds to SGD behavior near a saddle point. This approximation is also extensively utilized in the literature, particularly in trust-region methods [56, 57]. In this approach, the objective function is decomposed into a sequence of smaller subproblems, each of which is approximated by a quadratic model, thereby facilitating more efficient optimization by focusing on local regions of the solution space.
When $\sigma=0$, implying that each gradient is computed based on the entire training dataset such that the gradient noise is negligible, Eq. (13) reduces to the deterministic differential equation: $\mathrm{d}x_{i}(t)=-v_{i}x_{i}(t-\tau(t))\mathrm{d}t$. In addition, when $\tau(t)=0$, which corresponds to synchronous SGD, we obtain $\mathrm{d}x_{i}(t)=-v_{i}x_{i}(t)\mathrm{d}t+\sigma\mathrm{d}B(t)$, which corresponds to an OU process. Also, it is observed that the OU process with delay (11) along each eigenvector direction is independent, allowing us to first focus on the one-dimensional case (13) then extend our discussion to the multi-dimensional case.
Next, we demonstrate that the effects of the gradient noise and the gradient staleness can be separately considered taking the case that $\tau$ is a constant as an example. [^5] Note that the gradient staleness $\tau$ in SDDE is approximately a constant the variance of computation delay of each worker is small with a large worker number in ASGD, as will be demonstrated in Section IV. In this case, the analysis of the behavior of SGD can be based on viewing SGD as a discretization of the following associated stochastic delay differential equation, i.e.,
@@
Let $\tilde{x}_{d}(t)$ denote a solution to Eq. (15). Then, the solution to Eq. (14) can be given by $x_{s}(t)=\tilde{x}_{d}(t)+\sigma\int_{0}^{t}x_{0}(t-s)\mathrm{d}B(t),$ in which $x_{0}(t)=\sum_{k=0}^{\lfloor\frac{t}{\tau}\rfloor}\left[\frac{{v}^{k}}{k!}(t-k\tau)\right]$ in which $t\in(0,\infty)$. In addition, we have $\tilde{x}_{d}(t)=x_{0}(t)\varphi(0)+{v}\int_{-\tau}^{0}x_{0}(t-s-\tau)\varphi(s)\mathrm{d}s$. It can be noticed that the difference between $x_{s}(t)$ and $\tilde{x}_{d}(t)$ is only given by a Gaussian distributed random variable, the variance of which is jointly determined by the gradient staleness $\tau$ and the standard variance of the Wiener process $\sigma$.
In the following discussions, we will present performance analyses of the convergence rate of distributed SGD, considering the effects of gradient noise and gradient staleness, through the SDDE-based approach [58, 59].
Remark:  In this paper, we adopt the general assumption that the gradient noise follows a Gaussian distribution in Assumption 2. This Gaussian assumption is widely adopted in the literature, as demonstrated in works such as [34, 44, 60, 61]. However, the gradient noise can exhibit non-Gaussian characteristics, especially with small minibatch sizes where the central limit theorem fails. For example, the gradient noise is shown to be heavy-tailed through empirical evidence in [62]. In case the gradient noise is far from Gaussian distribution, the optimization path can be modeled as a delayed Lévy-driven OU process. More specifically, the optimization path can be characterized by the following SDDE with a non-Gaussian Lévy motion,
$\mathrm{d}x(t)=-{v}x(t-\tau)\mathrm{d}t+\mathrm{d}L\acute{e}(t),$ (16)
where $L\acute{e}(t)$ is a scalar Lévy motion and $L\acute{e}(0)=0$. More specifically, a scalar Lévy motion is a stochastic process with stationary and independent increments. That is, for any $s,t$ with $0\leq s<t$, the distribution of $L\acute{e}(t)-L\acute{e}(s)$ only depends on $t-s$, and for any $0\leq t_{0}<t_{1}<\cdots<t_{n}$, $L\acute{e}({t_{i}})-L\acute{e}(t_{i-1}),i=1,\cdots,n$, are independent.